OpenStudy (anonymous):

Use the Pythagorean Theorem to determine the distance between points A(5, 3) and B(-3, -3). 1) 64 2) 36 3) 100 4) 10

OpenStudy (anonymous):

Formula \[Distance= \sqrt{(x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}} \] insert yout values\[Distance= \sqrt{(5--3)^{2}+(3--3)^{2}} \] \[= \sqrt{8^{2}+6^{2}} \] \[= \sqrt{64+36} \] \[= \sqrt{100}=10 \]